Lucy and her bike together have a mass of 120 kg she slows down from 4.5m/s to 3.5 m/s how much kinetic energy does she lose


1
Expert's answer
2021-11-17T10:39:31-0500

We can find the losses in kinetic energy as follows:


"KE_{loss}=\Delta KE=KE_f-KE_i,""KE_{loss}=\dfrac{1}{2}m(v_f^2-v_i^2),""KE_{loss}=\dfrac{1}{2}\times120\ kg\times((3.5\ \dfrac{m}{s})^2-(4.5\ \dfrac{m}{s})^2)=-480\ J."